Ron sold 798 towles and soad despensers. Towles were sold at $6.00 each and despensers for $2.50 each. The total sales came to $3,223.50. How many of each were sold?
Answer by malakumar_kos@yahoo.com(315)


Ron sold 798 towles and soad despensers. Towles were sold at $6.00 each and despensers for $2.50 each. The total sales came to $3,223.50. How many of each were sold?

Let the number of towels be = x
and number of despensers be = y
total number of towels and despensers sold = 798
i.e, x+y = 798......(1)
cost of 1 towel = $6.00
cost of x towels = $6x
cost of 1 despensers = $2.50
cost of y despensers = $y(2.50)
total cost of x towels and y despensers = $3223.50
6x+y(2.50) = $3223.50........(2)
substituting for x from we get 6(798-y)+2.50y = 3223.50
4788-6y+2.50y = 3223.50
4788-3223.50 = 6y-2.50y
1554.50 = 3.50y
y = 1554.50/3.50= 444
x+y = 798
x = 798-444 = 354 .
therefore number of towels sold = 354
number of despensers sold = 444